MEGHÍVÓ

2004. január 26-án, hétfőn, a Nagyteremben 14 órakor kezdődő Összintézeti Szemináriumunk programja:

OLAV NJASTAD (Norvégia)

ORTHOGONALITY ON THE UNIT CIRCLE

című előadása.

The talk will give an overview of basic properties of orthogonal polynomials on the unit circle (Szegő polynomials) and related unctions, including discussion of some recent progress. Topics that will be dealt with are recursion formulas, properties of zeros, paraorthogonal polynomials and quadrature formulas, associated continued fractions and rational interpolation, trigonometric moment problems.

MINDEN ÉRDEKLŐDŐT SZERETETTEL VÁRUNK!

2004. január 26-án, hétfőn, a Nagyteremben 14 órakor kezdődő Összintézeti Szemináriumunk programja:

OLAV NJASTAD (Norvégia)

ORTHOGONALITY ON THE UNIT CIRCLE

című előadása.

The talk will give an overview of basic properties of orthogonal polynomials on the unit circle (Szegő polynomials) and related unctions, including discussion of some recent progress. Topics that will be dealt with are recursion formulas, properties of zeros, paraorthogonal polynomials and quadrature formulas, associated continued fractions and rational interpolation, trigonometric moment problems.

MINDEN ÉRDEKLŐDŐT SZERETETTEL VÁRUNK!